# A Dirichlet stochastic block model for composition-weighted networks

### v1

@@ Appendix B Classification EM with hybrid log-likelihood
The updates for the Dirichlet concentration parameter matrix $\mathbf{A}$ are not available in closed form and so are found numerically using the R function $optim$ (R Core Team,, 2019). L-BFGS-B optimisation procedure (Byrd et al.,, 1995) is used to update the set of parameters $\{\alpha_{kh}\}_{k,h=1}^{K}$ as it allows a permitted range of values to be set; in the case of DirSBM, we are only interested in strictly positive solutions.
@@ Appendix C Integrated completed likelihood
The second term of Equation (29) is derived in the same fashion as in Daudin et al., (2008), using a non-informative Jeffrey prior and Stirling formula for an approximation of a Gamma function, leading to
@@
Appendix D Simulation studies: additional notes
@@ E.2 K=5
$\mathbf{\hat{A}}=\begin{pmatrix}11.004&2.214&0.323&1.189&0.213\\ 12.197&4.057&0.847&1.984&0.282\\ 5.539&3.083&1.174&0.959&0.301\\ 3.035&1.298&0.231&0.158&0.105\\ 0.630&0.323&0.124&0.155&0.108\\ \end{pmatrix}.$ (39)
@@ Appendix F Bike sharing data: parameter estimates
Dirichlet parameter matrix $\mathbf{A}$ estimate of the 4-cluster DirSBM on the bike sharing data is

### v2

@@ Appendix B Classification EM with hybrid log-likelihood
The updates for the Dirichlet concentration parameter matrix $\mathbf{A}$ are not available in closed form and are found numerically using the R function $optim$ (R Core Team,, 2024). L-BFGS-B optimisation procedure (Byrd et al.,, 1995) is used to update the set of parameters $\{\alpha_{kh}\}_{k,h=1}^{K}$ as it allows a permitted range of values to be set; in the case of DirSBM, we are only interested in strictly positive solutions.
@@ Appendix C Integrated completed likelihood
The second term of Equation (28) is derived in the same fashion as in Daudin et al., (2008), using a non-informative Jeffreys prior and Stirling formula for an approximation of a gamma function, leading to
@@
Appendix D Simulation studies: additional details
@@ E.2 K=5
$\mathbf{\hat{A}}=\begin{pmatrix}11.004&2.214&0.323&1.189&0.213\\ 12.197&4.057&0.847&1.984&0.282\\ 5.539&3.083&1.174&0.959&0.301\\ 3.035&1.298&0.231&0.158&0.105\\ 0.630&0.323&0.124&0.155&0.108\\ \end{pmatrix}$ (38)
@@ Appendix F Bike sharing data: parameter estimates
The Dirichlet parameter matrix $\mathbf{A}$ estimate of the 4-cluster DirSBM on the bike sharing data is